OpenStudy (vera_ewing):

Given an exponential function for compounding interest, A(x) = P(.82)^x, what is the rate of change? A. –18% B. –8% C. 0.82% D. 82%

OpenStudy (vera_ewing):

@welshfella 82%?

OpenStudy (welshfella):

No - i think as the principal is decreasing Rate of change = 100-82 = -18% Not absolutely sure though

OpenStudy (vera_ewing):

@sdfgsdfgs It's -18?

OpenStudy (welshfella):

yes you'd better check

OpenStudy (amoodarya):

\[a(x)=p*(1+r)^n\]

OpenStudy (sdfgsdfgs):

the prob is not clear: "rate of change" is not uniform for an exponential function... http://www.regentsprep.org/regents/math/algebra/AE7/ExpDecayL.htm if it is asking for "growth rate", then @welshfella is right on -18%.

OpenStudy (vera_ewing):

And if it's not asking for "growth rate" then would the answer be different?

OpenStudy (sdfgsdfgs):

then there be no ans :P

OpenStudy (vera_ewing):

Okay. Thanks!

OpenStudy (sdfgsdfgs):

welcome